EMISSION SPECTROSCOPY AND AB INITIO CALCULATIONS ON VCl AND VF

\underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; J. LI’EVIN, Université Libre de Bruxelles, Laboratoire de Chimie Physique Moléculaire, CP 160/09, Av. F. D. Roosevelt 50, Bruxelles, Belgium; P. F. BERNATH, Department of Chemistry, University of Waterloo, Ont., Canada N2L 3G1; S. P. DAVIS, Department of Physics, University of California, Berkeley, CA 94720.